# Unveiling the Potential of Multimodal Retrieval Augmented Generation with Planning

### v1

@@ 6.2.1. End-to-end Performance
Table 2 presents a comprehensive comparison of the end-to-end performance comparison of various MLLMs integrated with current MRAG methodologies and our proposed CogPlanner. The following observations can be drawn based on these results: (1) Enhanced Performance with CogPlanner: Notably, CogPlanner with GPT-4o consistently yields best performance compared to all other configurations. Specifically, CogPlanner results in a significant improvement over MRAG baselines, with performance gains ranging from 12.4% to 52.5%, as measured by the average end-to-end performance metrics. This enhancement is attributed to its ability to decompose and refine complex, multimodal queries. By simplifying these queries, CogPlanner facilitates the dynamic determination of necessary retrieval actions, thereby ensuring the acquisition of accurate, complementary information. These results highlight the critical role of MRAG Planning in optimizing the performance of MRAG systems. (2) Weakness of Fixed Search Strategies: The two fixed search strategies, while offering some improvements in specific metrics, generally exhibit a negative impact when compared to direct responses generated by MLLMs. As anticipated, these rigid search actions, particularly when applied to concise queries and screenshot images within multimodal queries, introduce noise that misleads downstream MLLMs rather than providing useful information. The noisy retrieval results tend to obscure the relevant information, thereby diminishing the effectiveness of the system. These findings further validate the motivation behind the development of the CogPlanner framework. (3) Comparison of Parallel and Sequential Modeling Approaches: Comparing our parallel and sequential modeling methodologies, they show comparable performance, each demonstrating significant improvements over traditional MRAG systems. However, the sequential modeling approach does not exhibit substantial advantage over the parallel approach. This suggests that the current capabilities of MLLMs are insufficient for accurately evaluating complementary information in a sequential manner (4) Benefits on Multihop Query: The performance of CogPlanner shows variability across 1-3 hop query categories, with more pronounced improvements observed in multi-hop queries. This pattern suggests that the iterative and adaptive planning process of CogPlanner is particularly beneficial in scenarios that require the retrieval of information from multiple sources and involve multi-step reasoning to formulate a complete and accurate response.
Table 4. End-to-end performance comparison of Qwen2-VL-72B-Instruct and Qwen2-7b-VL-Cog as planning experts, with response generation models consistently using Qwen2-VL-72B-Instruct.
Planning Expert Precision Recall F1 Avg
CogPlanner With Parallel Modeling
Qwen2-VL-72B-Instruct 0.3245 0.3690 0.2174 0.3036
Qwen2-7b-VL-Cog 0.3197 0.3265 0.2146 0.2869
CogPlanner With Sequential Modeling
Qwen2-VL-72B-Instruct 0.3179 0.3633 0.2065 0.2959
Qwen2-7b-VL-Cog 0.3250 0.3310 0.2138 0.2900
@@ 6.2.2. Query Reformulation Performance
Table 5. The efficiency evaluation of CogPlanner overhead in MRAG systems, with the ratio calculated based on the average token count of the generated response for each experimental setting.
Planning Expert Category # Para Tokens # Total Tokens
Qwen2-VL-72B-Instruct Parallel 8.90 (13.7%) 9.76 (14.9%)
Sequential 13.56 (21.6%) 13.56 (21.6%)
Qwen2-7b-VL-Cog Parallel 6.59 (8.5%) 7.58 (9.8%)
Sequential 7.59 (11.9%) 7.59 (11.9%)
@@ 6.3. Lightweight Model Performance
In this section, we assess the performance of our fine-tuned Qwen2-7b-VL-Cog from both effectiveness and efficiency perspectives.
@@ 6.3.1. Model Evaluation
To ensure a fair comparison, we use the Qwen2-VL-72B-Instruct as the reference model for final answer generation. Both Qwen2-VL-72B-Instruct and our fine-tuned Qwen2-7b-VL-Cog are utilized as planning experts within the CogPlanner framework, facilitating an assessment of whether smaller MLLMs, specifically through our tailored supervised fine-tuning procedure, can effectively manage the complexities inherent in MRAG planning tasks. The results, as presented in Table 4, underscore the critical role of model selection in determining the overall performance. While larger MLLMs typically demonstrate superior results, our Qwen2-7b-VL-Cog closely approximates the performance of Qwen2-VL-72B-Instruct across most evaluation metrics. This finding serves to validate the efficacy of our fine-tuning strategy. Under the CogPlanner framework, our results demonstrate that it is indeed possible to deploy a resource-efficient planning expert and lead to enhanced performance of the MRAG system. Specifically, the Qwen2-7b-VL-Cog emerges as a compelling alternative.
@@ 6.3.2. Efficiency Evaluation
We conduct a comparison of the efficiency of CogPlanner between Qwen2-VL-72B-Instruct and Qwen2-7b-VL-Cog within the context of MRAG systems, focusing on the overall cost of the planning procedure. We utilize token generation as the primary metric for evaluating efficiency. Specifically, the ”Total Tokens” refers to the cumulative number of tokens generated by the MLLMs during the entire planning process. Additionally, we introduce the ”Para Token” metric as a near-latency measure. Given that the parallel modeling paradigm uses two threads for decision-making across two actions, we consider the maximum token count generated during each round of planning as the cost for that round. To provide a more comprehensive understanding of the additional computational cost incurred by incorporating CogPlanner, we calculate the average number of tokens in the generated response and compute the ratio between the tokens used in planning versus those needed for answering. The evaluation results are summarized in Table 5. Our findings indicate that Qwen2-7b-VL-Cog is a significantly more lightweight module for MRAG systems compared to Qwen2-VL-72B-Instruct, introducing only a 10% increase in computational cost. This result underscores the minimal additional overhead imposed by the lightweight integration of CogPlanner into MRAG systems, highlighting its suitability for real-world applications. Furthermore, the parallel execution model exhibits superior performance relative to the sequential modeling approach across both efficiency metrics, aligning with the design expectations.

### v2

@@ 6.2.1. End-to-end Performance
Table 2 presents a comprehensive comparison of the end-to-end performance of various MLLMs integrated with current MRAG methodologies and our proposed CogPlanner. The following observations can be drawn based on these results: (1) Enhanced Performance with CogPlanner: Notably, CogPlanner with GPT-4o consistently yields best performance compared to all other configurations. Specifically, it delivers substantial improvements over baseline MRAG systems, with end-to-end performance gains ranging from 12.4% to 52.5%, and at least a 41.45% improvement over self-reflective MRAG variants. This enhancement is attributed to its ability to decompose and refine complex queries. By simplifying these queries, CogPlanner facilitates the dynamic determination of necessary retrieval actions, thereby ensuring the acquisition of accurate, complementary information. These results highlight the critical role of MRAG Planning in optimizing the performance of MRAG systems. (2) Weakness of Fixed Search Strategies: The two fixed search strategies, while offering some improvements in specific metrics, generally exhibit a negative impact when compared to direct responses generated by MLLMs. As anticipated, these rigid search actions, particularly when applied to concise queries and screenshot images within multimodal queries, introduce noise that misleads downstream MLLMs rather than providing useful information. The noisy retrieval results tend to obscure the relevant information, thereby diminishing the effectiveness of the system. These findings further validate the motivation behind the development of the CogPlanner framework. (3) Comparison of Parallel and Sequential Modeling Approaches: Comparing our parallel and sequential modeling methodologies, they show comparable performance, each demonstrating significant improvements over traditional MRAG systems. However, the sequential modeling approach does not exhibit a substantial advantage over the parallel approach. This suggests that the current capabilities of MLLMs are insufficient for accurately evaluating complementary information in a sequential manner (4) Benefits on Multihop Query: The performance of CogPlanner shows variability across 1-3 hop query categories, with more pronounced improvements observed in multi-hop queries. This pattern suggests that the iterative and adaptive planning process of CogPlanner is particularly beneficial in scenarios that require the retrieval of information from multiple sources and involve multi-step reasoning to formulate a complete and accurate response.
@@ 6.3. Lightweight Model Performance
In this section, we assess the performance of our fine-tuned Qwen2-7B-VL-Cog from both effectiveness and efficiency perspectives. The results are presented in Table 4.
@@ 6.3.1. Model Evaluation
To ensure a fair comparison, we use the Qwen2-VL-72B-Instruct as the reference model for final answer generation. Both Qwen2-VL-72B-Instruct and our fine-tuned Qwen2-7B-VL-Cog are utilized as planning experts within the CogPlanner framework, facilitating an assessment of whether smaller MLLMs, specifically through our tailored supervised fine-tuning procedure, can effectively manage the complexities inherent in MRAG planning tasks. The results underscore the critical role of model selection in determining the overall performance. While larger MLLMs typically demonstrate superior results, our Qwen2-7B-VL-Cog closely approximates the performance of Qwen2-VL-72B-Instruct across most evaluation metrics. This finding serves to validate the efficacy of our fine-tuning strategy. Under the CogPlanner framework, our results demonstrate that it is indeed possible to deploy a resource-efficient planning expert, leading to enhanced performance of the MRAG system. Specifically, the Qwen2-7B-VL-Cog emerges as a compelling alternative.
@@ 6.3.2. Efficiency Evaluation
We compare the efficiency of CogPlanner with Qwen2-VL-72B-Instruct and Qwen2-7B-VL-Cog as the planning expert. Specifically, we use total token generation and latency as the primary evaluation metrics, excluding the cost associated with final answer generation to isolate planning efficiency. As shown in Table 4, our findings indicate that Qwen2-7B-VL-Cog constitutes a significantly more lightweight module for MRAG systems, incurring only a 10% increase in token consumption, but reducing latency to just 30% compared to the Qwen2-VL-72B-Instruct model. It emerges as a practical compromise between performance and computational efficiency, particularly well-suited for deployment in real-world industrial settings. Furthermore, the parallel execution model exhibits superior performance relative to the sequential modeling approach across both efficiency metrics, aligning with our design expectations.